x+1/x=5, then find the value of x^16+1/x^16 ?


1
Expert's answer
2021-01-29T04:44:58-0500

Given

"x+1\/x=5"

Tahing square on both sides we have

"(x+1\/x)^2=5^2"

solving the bracket by using formula "(a+b)^2=a^2+2ab+b^2"

we get

"x^2+2(x)(1\/x)+1\/x^2=25,"

"x^2+1\/x^2=25-2"

"=x^2+1\/x^2=23"

again taking aquare on both sides we have ,

"(x^2+1\/x^2)^2=23^2"

"=x^4+(2)(x^2)(1\/x^2)+1\/x^4=529,"

"x^4+1\/x^4=529-2"


"=x^4+1\/x^4=527"


taking aquare on both sides we have

"=(x^4+1\/x^4)^2=(527)^2"

"=x^8+(2)(x^4)(1\/x^4)+1\/x^8=277729,"

"x^8+(2)+1\/x^8=277729-2"

"=x^8+1\/x^8=277727"

taking aquare on both sides we have

"=(x^8+1\/x^8)^2=(277727)^2"

"=x^{16}+(2)(x^8)(1\/x^8)+1\/x^{16}=(277727)^2,"

"x^{16}+1\/x^{16}=(277727)^2-2"





Answer:

"x^{16}+1\/x^{16}=(277727)^2-2"